OpenStudy (anonymous):

find the distance between the points (3,5) and (1,-8)

OpenStudy (anonymous):

Use the distance formula! c: d= [(SQUARE ROOT OVER THE WHOLE THING BTW NOT JUST (X1-X2)^2) (x1-x2)@2 + (y1-y2)^2]

OpenStudy (anonymous):

so d= SQUARE ROOT* (-1-5)^2 + (3-0)^2 d= square root* (-6)^2 + (3)^2 d= square root* 36+9 d=square root of 45 d= 3 root 5